# Heterogeneous Intertemporal Treatment Effects via Dynamic Panel Data Models

### v1

@@ 4 Evaluating Sequential Exchangeability
In this section, we assess the plausibility of sequential exchangeability (SE) in Assumption 2, which has played a central role in our results and discussion in Section 3. This condition amounts to a form of “selection on observables”, as it posits (mean or full) independence between potential outcomes and contemporary treatment conditional on observed history of past outcomes and treatments.
Our objective is to provide guidelines for empirical researchers to decide when and how such restrictions on selection-on-observables can be justified in observational or experimental data. Naturally, justifying selection on observables in this dynamic setting would require restrictions on how the sequence of treatments are determined.
In Section 4.1 and 4.2, we showcase the nature of these restrictions within a structural model where treatments are chosen by rational, forward-looking individual units with learning, and compare them with the parallel trends condition in the DiD literature. In Section 4.3, we show that sequential exchangeability holds in an experimental design of sequentially randomized treatment (where the treatment in period $t$ is randomized conditioning on the history of past treatments and outcomes); in contrast, the parallel trends conditions in general would not hold in such a case.
4.1 A Structural Model of Treatment Choices with Learning
We focus on a structural model of dynamic treatment choices that enables individual decision-makers to be both forward-looking and capable of learning. Learning is an interesting possibility in dynamic choice which is potentially compatible with sequential exchangeability assumptions because of the latter’s dependence on the history of observed outcomes. Unlike Section 3.4, the model studied in this section refrains from any parametric specification for the potential outcomes; instead, the model emphasizes scenarios in which forward-looking individuals make rational treatment decisions, continually learning from previous treatments and outcomes. To simplify the notation in this section, we omit the individual subscripts $i$.
In each period $1\leq t\leq T$, the potential outcome is determined by:
$Y_{t}(d^{t})=\lambda_{t}(d^{t},Y^{t-1}(d^{t-1}),\alpha(d_{t}),\varepsilon_{t}(d_{t}))\text{,}$ (27)
where $Y^{s}(d^{s})=(Y_{0},Y_{1}(d^{1}),\dots Y_{s}(d^{s}))$ denotes a vector of counterfactual outcomes, with convention $Y^{0}(d^{0})=Y_{0}$. The initial condition $Y_{0}$ is allowed to be correlated with subsequent potential outcomes $Y_{t}(d^{t})$ for $t\geq 1$. We refer to $\alpha(\cdot)\equiv(\alpha(1),\alpha(0))$ as (stochastic) unobserved individual heterogeneity, or fixed effects, and $\varepsilon_{t}(\cdot)$ as potential, transitory noises affecting potential outcomes at time $t$. Individual decision makers do not observe $\alpha(\cdot)$ or $\varepsilon_{t}(\cdot)$, but know the deterministic function $\lambda_{t}(\cdot)$ as well as the distribution of $\varepsilon_{t}(\cdot)$ for all $t$555 Note that the structural dependence on the lagged outcome generalizes indexing of the full decision history because the outcome-generating function $\lambda_{t}(\cdot)$ is deterministic and depends only on present noise $\varepsilon_{t}(d_{t})$; in particular, the lagged outcome permits some dependence on the past noise vector $\varepsilon^{t-1}(d^{t-1})$.
Each individual holds a prior belief for $\alpha(\cdot)$ at $t=1$, and uses incremental information to update their beliefs for $\alpha(\cdot)$ at the beginning of each period $t\geq 2$. Specifically, the individual’s prior over $\alpha(\cdot)$ at the start of $t=1$ is $\Psi_{1}(\cdot;\psi_{1})$, which is fully characterized by a random vector $\psi_{1}$. Let $\psi_{1}\equiv\varphi_{1}(H_{0},\xi_{0})$, where $H_{0}=Y_{0}$ denotes the observed history and $\xi_{0}$the unobserved individual factors in period 0 that determine the initial prior in period 1. In period $t\geq 2$, the posterior for $\alpha(\cdot)$ is $\Psi_{t}(\cdot;\psi_{t})$, which is completely characterized by an updated vector $\psi_{t}=\varphi_{t}(\psi_{t-1},H_{t-1},\xi_{t-1})$, where $H_{t}\equiv(D_{t},Y_{t})$ is the incremental information (observed by the researcher) that affects the individual’s posterior, and $\xi_{t}(d^{t})$ is the incremental information (unobserved by the researcher) from choice $d_{t}$ in period $t$, with $\xi_{t}=\xi_{t}(D^{t})$. Let $H^{t}\equiv(H_{0},\dots,H_{t})$ and $\xi^{t}=(\xi_{0},\dots,\xi_{t})$ denote the realized vectors and $\xi^{t}(\cdot)$ the full potential vector through period $t$. Using recursive substitution and with a slight abuse of notation, we can write: [^6] With $\Psi_{t}$ and $\varphi_{t}$ indexed by $t$, we allow for non-conjugate beliefs. Also, $\varphi_{t}$ may depend on unknown parameters, such as the variance of $\varepsilon_{t}(\cdot)$ in a normal learning model.
$\psi_{t}=\varphi_{t}(H^{t-1},\xi^{t-1}).$
To reiterate, we assume that $\psi_{t}$ is a sufficient statistic for the posterior of $\alpha(\cdot)$ at time $t$ given the outcome-relevant history $(H^{t-1},\xi^{t-1})$ observed by the decision-maker.
Let $\mathcal{D}^{T}$ denote the set of possible sequences of choices; note that, in contrast to preceding sections, we do not assume that choices in each period are binary. Given a choice history $D^{t-1}=d^{t-1}$, an individual’s instantaneous, ex post utility from choosing $d_{t}$ at time $t$ is:
$u_{t}\left(d_{t},Y_{t}(d^{t}),X_{t}\right)+\eta_{t}(d_{t})$
where $X_{t}$ are additional states (apart from outcomes $Y_{t}$), and $\eta_{t}(\cdot)$ are potential shocks. Individuals observe $(X_{t},\eta_{t}(\cdot))$ and know the deterministic function $u_{t}(\cdot)$ for all $t\geq 1$. We consider the case where $X_{t}$ consists of persistent or serially correlated heterogeneities that are not reported in the data; conceptually, it is not hard to extend to the cases where some components in $X_{t}$are also reported in the sample. To allow the transition of states to depend on choice, it is useful to also define the potential state $X_{t}(d^{t-1})$ that would have occurred in period $t$ under choice history $d^{t-1}$; as usual, $X^{T}=X^{T}(D^{T-1})$.
Summarizing, in each period $t$, the individual observes $(\eta_{t}(\cdot),X_{t})$ before choosing $D_{t}$, and observes $(H_{t},\xi_{t})$ after choosing $D_{t}$. The individual knows the full joint distribution of potential random variables across periods, but not the realizations except as previously specified. For example, the individual does not generally observe the realization of $\alpha(\cdot)$ or $\varepsilon_{t}(\cdot)$ in the outcome-generating process. Define $\zeta^{t}(d^{t-1})\equiv(\xi^{t-1}(d^{t-1}),X^{t}(d^{t-1}),\eta^{t}(\cdot))$ as the potential accumulation of the individual’s information in period $t$ given potential choice history $d^{t-1}$, which is not reported in the data. With minor loss of generality, we consider choices $D_{t}$ that are a function of the individual’s observed and unobserved information:
$D_{t}=\phi_{t}(H^{t-1},\zeta^{t})$ (28)
Unlike time-homogeneous, Markovian choice rules in a typical dynamic discrete choice model, e.g., Rust (1987), the (rational) rule of treatment choices in this section needs to be neither Markovian nor time-homogeneous. This is because the individual unit needs to update its beliefs in each period using the full history of outcomes and endogenous choices in the past.
The key sufficient condition for sequential exchangeability is that the unobserved information for any counterfactual choice sequence is essentially uninformative about the outcome process, conditional on the initial observed outcome:
$\zeta^{T}(d^{T-1})\perp(\alpha(\cdot),\varepsilon^{T}(d^{T}))\mid Y_{0}\quad\text{for all $d^{T}\in\mathcal{D}^{T}$.}$ (29)
For clarity at the expense of some generality, this nests several more primitive assumptions based on the unobserved vector $\zeta^{t}=(\xi^{t-1},X^{t},\eta^{t}(\cdot))$. First, there is no meaningful unobserved heterogeneity in prior beliefs through $\xi_{0}$, though prior beliefs may vary systematically with $Y_{0}$. Second, there is no meaningful unobserved heterogeneity in incremental learning through $\xi_{t}$ for $t\geq 1$, though the belief parameter $\psi_{t}$ may evolve systematically based on the observed history $H^{t-1}$. Third, the state transition process $X^{T}$ is uninformative about the outcome process, so that there is no learning based on the unobserved evolution of states. Finally, independence of the shocks in instantaneous ex post utilities $\eta_{t}(\cdot)$ is standard. In summary, (29) permits systematic learning based only on observables. In that case, (mean) sequential exchangeability is satisfied.
Proposition 6
Suppose (29) holds in the structural choice model with learning. Then sequential exchangeability (Assumption 2) is satisfied.
The intuition underlying this result is straightforward. By definition (27), for any sequence of potential treatment $d^{t}$, the history of potential outcomes $Y^{t}(d^{t})$ is a function of initial condition $Y_{0}$, fixed effects $\alpha(\cdot)$ and noises $\varepsilon^{t}(d^{t})$ only. Meanwhile, conditional on an observed history $D^{s-1}=d^{s=1}$, $Y^{s-1}(d^{s-1})=y^{s-1}$, any event $D_{s}=d_{s}$ is defined over the sample space of choice unobservables $\zeta^{s}(d^{s-1})$, which is conditionally independent from the outcome unobservables, and consequently the sequence of potential outcomes. This means the confounders for the effect of $D_{s}$ on subsequent potential outcomes $Y_{t}(d^{s},\cdot)$ are also controlled for in this case under sequential exchangeability.
Our interpretation of this result is twofold. First, it highlights the limitations of sequential exchangeability in a dynamic choice context. In this sense, the result is more about specifying the channels of unobserved choice heterogeneity that must be ruled out, than about what is allowed. Thus, even as the choice model recovers the possibility of fixed effects in the outcome model without violating sequential exchangeability, it imposes significant restrictions on unobserved heterogeneity in choice.
At the same time, viewed in a more positive light, Proposition 6 provides sufficient conditions under which sequential exchangeability can be used to identify the causal effects when treatment choices are made by forward-looking individuals capable of learning. In this case, sequential exchangeability provides an alternative suite of methods for recovering treatment effects in the presence of learning under (limited) unobserved heterogeneity, complementary to the pure learning model of Bunting, Diegert, and Maurel (2022).
4.2 Comparison With Parallel Trends: Under Learning
We next compare sequential exchangeability and parallel trends within the context of the structural learning model in Section 4.1. It is most illustrative to make such a comparison in a simplified version of the model with no additional states $X_{t}(d^{t-1})$ and $T=2$. To reiterate, in such a simplified case, the potential outcomes are:
$Y_{1}(d_{1})$ $=$ $\lambda_{1}(d_{1},Y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}))\text{;}$ (30)
$Y_{2}(d_{1},d_{2})$ $=$ $\lambda_{2}(d_{1},d_{2},Y_{0},Y_{1}(d_{1}),\alpha(d_{2}),\varepsilon_{2}(d_{2}))\text{,}$
and the rules of treatment choices are:
$D_{1}$ $=$ $\phi_{1}(Y_{0},\xi_{0},\eta_{1}(\cdot))\text{;}$ (31)
$D_{2}$ $=$ $\phi_{2}(Y_{0},Y_{1},D_{1},\xi^{1}(D_{1}),\eta_{2}(\cdot))\text{,}$
where $\xi^{1}(d_{1})\equiv(\xi_{0},\xi_{1}(d_{1}))$. [^7] Recall $D_{0}=0$ w.p.1 and is therefore dropped from $\xi_{0}(D_{0}),\xi_{1}(D_{0},D_{1})$ for simplicity.
In the current setting with dynamic treatment effects, the parallel trends assumption (conditional on initial outcomes $Y_{0}$) requires:
@@ 4.2 Comparison With Parallel Trends: Under Learning
See, e.g., de Chaisemartin and D’Haultfoeuille (2024). The next proposition shows, under a conditional independence condition, sequential exchangeability holds whereas parallel trends would not hold in general.
Proposition 7
Consider the model Section 4.1 with no states $X_{t}(d^{t-1})$ and $T=2$. If
$(\alpha(\cdot),\varepsilon^{2}(d^{2}))\perp(\xi^{1}(d_{1}),\eta^{2}(\cdot))\mid Y_{0}\text{ for all }d^{2}\equiv(d_{1},d_{2})\text{,}$ (33)
then sequential exchangeability in Assumption 2b holds while $E[Y_{2}(0,0)-Y_{1}(0)|D_{1},D_{2},Y_{0}]$ is in general a non-degenerate function of $(D_{1},D_{2})$.
The conditional independence in (33) specializes (29) when there are no additional states $X_{t}(d^{t-1})$ and $T=2$. This condition posits the unobserved states influencing posterior formation are independent from the individual fixed effects, after controlling for the initial outcome $Y_{0}$; it does not further restrict the learning process.
For example, consider a simplistic case where $\varepsilon^{2}(d^{2}),\eta^{2}(\cdot)$ are idiosyncratic noises independent from all other variables, and where $\xi^{t}(d^{t})=\xi_{0}$. That is, the evolution of posterior beliefs only depends on an initial shifter of the prior $\xi_{0}$ and observed history $(D^{t-1},Y^{t-1})$, but no incremental, time-varying $\xi_{s}(d_{s})$ for $s\geq 1$. Then (33) holds when $\alpha(\cdot)$ is independent from the initial prior shifter $\xi_{0}$ given $Y_{0}$.
The result in Proposition 7 is intuitive. With forward-looking and learning, the treatment choices $D^{2}=(D_{1},D_{2})$ are a non-trivial function of a realized history of observed outcomes $Y_{0},Y_{1}(D_{1})$. Hence conditioning on different realizations of $D^{2}$ amounts to conditioning on different events defined over the sample space of $(Y_{0},\alpha(\cdot),\varepsilon_{1}(\cdot),\xi^{1}(\cdot),\eta^{2}(\cdot))$. Because $Y_{2}(0,0)-Y_{1}(0)$ is a function of $(Y_{0},\alpha(\cdot),\varepsilon^{2}(\cdot))$, its distribution (and mean) generally vary with the conditioning event defined by $D^{2}$. This non-degeneracy holds in general, even after controlling for the realization in $Y_{0}$.
@@ 4.3 Sequentially Randomized Treatments
We conclude this section with yet another intuitive and important case of treatment assignment mechanism, in which sequential exchangeability holds but parallel trends do not. This is the case where the treatment assignments are sequentially randomized conditional on earlier history of treatments and outcomes.
Suppose the treatments are sequentially randomized as follows:
@@
Such a treatment assignment mechanism is relevant especially in experimental designs where a researcher may be able to assign randomized treatment in each period after controlling for observed outcomes up to that period.
For simplicity, let the initial outcome $Y_{0}$ be degenerate at a constant value and suppress it in all conditioning events below. Mean sequential exchangeability in Assumption 2b holds under (34) and (35), because
@@
In contrast, the parallel trends condition in (32) would fail in general under (34) and (35). To see this, note:
@@
where $\mu_{2}(s)\equiv E[Y_{2}(0,0)|Y_{1}(0)=s,D_{1}=0]$ is a function of $s$ that does not vary with $d_{2}$, and the second equality is due to (35). Besides, by construction,
[…]
@@
where $p(s)\equiv\Pr\{D_{2}=1|Y_{1}(0)=s,D_{1}=0\}$ and $G(s)\equiv\Pr\{Y_{1}(0)\leq s|D_{1}=0\}=\Pr\{Y_{1}(0)\leq s\}$. Thus, when $p(\cdot)$ is non-degenerate and non-linear, $F_{Y_{1}(0)}(y|D_{2}=d_{2},D_{1}=0)$ generally varies with $d_{2}$. As a result, (32) would not hold in general.
@@
Note that (32) holds in a special case where $p(\cdot)$ is degenerate, e.g., $D_{2}$ is randomly assigned independent of past $Y_{1}(\cdot)$. In such a case, $F_{Y_{1}(0)}(\cdot|D_{2}=d_{2},D_{1}=0)=G(\cdot)$ does not vary with $d_{2}$ and (32) holds.

### v2

@@ 4 Evaluating Sequential Exchangeability
So far, we have provided a causal interpretation of a first-differenced IV estimand under the sequential exchangeability (SE) condition in Section 3.2, and investigated two semiparametric structural models for which causal parameters, such as ATFT, are identified without SE in Section 3.3.
Our work complements the popular difference-in-differences method, which estimates ATT under a parallel trends (PT) assumption. In the two-period case of our setting with heterogeneous intertemporal treatment effects, the PT assumption, conditional on initial outcomes $Y_{0}$, amounts to:
@@
See, e.g., de Chaisemartin and D’Haultfoeuille (2024).
In this section, we study which kinds of relations between treatment selection and potential outcomes are permitted or ruled out under SE and PT respectively. This provides a framework for empirical researchers to assess the circumstances under which SE and PT are justified in experimental or observational data. Since we no longer distinguish between homogeneous and heterogeneous individual-level parameters as in Subsection 3.3, we again suppress indexing by individual effects.
In Section 4.1, we confirm that SE holds in an experimental design of sequentially randomized treatment (where the treatment in period $t$ is randomized conditioning on the history of past treatments and outcomes); in contrast (and perhaps surprisingly), PT would not generally hold in such a case.
In Section 4.2, we turn to observational data and showcase the nature of SE restrictions within a structural model where treatments are chosen by rational, forward-looking individuals with learning; again we compare these SE restrictions with the PT conditions.
@@ 4.1 Sequentially Randomized Treatments
Suppose the treatments are sequentially randomized conditional on earlier history of realized treatments and outcomes. For $T=2$, this means:
@@
Such a treatment assignment mechanism is relevant especially in experimental designs where a researcher may be able to assign randomized treatments in each period after controlling for observed outcomes up to that period.
We show that while the sequential randomization of treatments implies sequential exchangeability, it is generally not sufficient for parallel trends. For simplicity, let the initial condition $Y_{0}$ be degenerate (constant), and suppress it in all conditioning events.
Proposition 7
Sequential randomization of treatments in (29) and (30) implies the (mean) sequential exchangeability in Assumption 2b, but not the parallel trends in (28).
The proof of Proposition 7 is intuitive and included in the text. Mean sequential exchangeability in Assumption 2b holds under (29) and (30) because:
@@
In contrast, the parallel trends condition in (28) is not implied under (29) and (30). To see this, note:
@@
where $\mu_{2}(d_{2},s)\equiv E[Y_{2}(0,0)|D_{2}=d_{2},Y_{1}(0)=s,D_{1}=0]$ does not vary with $d_{2}$ under (30), and hence the first argument is suppressed. Besides, by construction,
$F_{Y_{1}(0)}(y|D_{2}=1,D_{1}=0)$ $=$ $\frac{\int_{-\infty}^{y}p(s)dG(s)}{\int_{-\infty}^{+\infty}p(s)dG(s)}\text{,}$
$F_{Y_{1}(0)}(y|D_{2}=0,D_{1}=0)$ $=$ $\frac{\int_{-\infty}^{y}[1-p(s)]dG(s)}{\int_{-\infty}^{+\infty}[1-p(s)]dG(s)}\text{,}$
@@
where $p(s)\equiv\Pr\{D_{2}=1|Y_{1}(0)=s,D_{1}=0\}$, and $G(s)\equiv\Pr\{Y_{1}(0)\leq s|D_{1}=0\}=\Pr\{Y_{1}(0)\leq s\}$ under (29). Thus, when $p(\cdot)$ is non-degenerate and non-linear, $F_{Y_{1}(0)}(y|D_{2}=d_{2},D_{1}=0)$ generally varies with $d_{2}$. Therefore, the parallel trends condition in (28) does not hold in general.
@@
Note that (28) does hold in a special case where $p(\cdot)$ is degenerate, e.g., $D_{2}$ is randomly assigned independent of past $Y_{1}(\cdot)$ conditional on $D_{1}=0$. In such a case, $F_{Y_{1}(0)}(\cdot|D_{2}=d_{2},D_{1}=0)=G(\cdot)$ does not vary with $d_{2}$ and (28) holds. This is consistent with results in Marx, Tamer, and Tang (2024) on failure of parallel trends in models where treatments are functions of past outcomes.
4.2 Forward-Looking Treatment Choices with Learning
We now study when the SE and PT conditions are justified in an observational setting, where individuals self-select into treatments through rational, forward-looking choices based on beliefs updated through learning.
Consider a structural model with $T=2$, where the potential outcomes are:
$Y_{1}(d_{1})$ $=$ $\lambda_{1}(d_{1},Y_{0},\alpha(d_{1}),\varepsilon_{1})\text{,}$ (31)
$Y_{2}(d_{1},d_{2})$ $=$ $\lambda_{2}(d_{1},d_{2},Y_{0},Y_{1}(d_{1}),\alpha(d_{2}),\varepsilon_{2})\text{,}$
with $\lambda_{t}(\cdot)$ being deterministic functions, $\alpha(\cdot)$ potential individual fixed effects that vary with contemporary counterfactual treatments $d_{t}$, and $\varepsilon_{t}$ idiosyncratic, time-varying structural errors in potential outcomes. The DGP for realized treatment choices are:
$D_{1}$ $=$ $\phi_{1}(Y_{0},\xi_{0},\eta_{1}),$ (32)
$D_{2}$ $=$ $\phi_{2}(Y^{1},D_{1},\xi_{1}(D_{1}),\eta_{2}),$
where $Y^{1}\equiv(Y_{0},Y_{1})\equiv(Y_{0},Y_{1}(D_{1}))$, $\phi_{t}(\cdot)$ are deterministic functions, and $\xi_{0}$ and $\xi_{1}(d_{1})$ are potential individual fixed effects which may evolve over time, with $\xi_{1}(d_{1})$ possibly subsuming $\xi_{0}$ as a subvector, and $\eta_{t}$ are idiosyncratic, time-varying noises. [^10] Recall that $D_{0}=0$ w.p.1, and is dropped from $\xi_{0}(D_{0}),\xi_{1}(D_{0},D_{1})$ for simplicity.
Our results below generalize to the cases with $T\geq 3$, or where the idiosyncratic noises $\varepsilon_{t}$ and $\eta_{t}$ are also indexed by contemporary counterfactual treatments $d_{t}$. Such generalizations do not require qualitatively different argument, and are omitted for conciseness.
We provide an example of a structural model where the treatment rule in (32) arises endogenously from the rational choices by forward-looking individuals with learning.
Example 1. (Dynamic treatment choices with learning.) Consider a model with two periods $T=2$. Each individual’s per-period, ex post payoff from choosing $D_{1}=d_{1}$ at $t=1$ is:
$u_{1}(d_{1},Y_{1}(d_{1}))+\eta_{1}(d_{1}).$
Moreover, given $D_{1}$, the instantaneous, ex post payoff from choosing $D_{2}=d_{2}$ at $t=2$ is:
$u_{2}(d_{2},Y_{2}(D_{1},d_{2}))+\eta_{2}(d_{2}).$
Potential outcomes $Y_{t}(\cdot)$ are determined by $(Y_{0},\varepsilon_{1},\varepsilon_{2})$ and individual fixed effects $\alpha\equiv(\alpha(0),\alpha(1))$ via deterministic functions $\lambda_{t}(\cdot)$ as defined in (31).
Individuals know all functions $u_{t}(\cdot)$ and $\lambda_{t}(\cdot)$, but never observe the potential fixed effects $\alpha$ and idiosyncratic errors $\varepsilon_{t}$ in $Y_{t}(\cdot)$. In each period $t$, individuals observe $\eta_{1}\equiv(\eta_{1}(0),\eta_{1}(1))$, and know the distribution of the idiosyncratic errors $\varepsilon_{t}$.
At the start of the first period $t=1$, each individual holds a prior belief for $\alpha$. This prior is characterized by a parameter vector $\kappa_{1}\equiv\varphi_{1}(Y_{0},\xi^{*})$, where $\xi^{*}$ is a persistent fixed effect known privately to the individual. Similarly, at the start of the next period $t=2$, the individual uses past observed outcomes $Y^{1}$ and the realized treatment $D_{1}$ to update the posterior for $\alpha$, which is characterized by $\kappa_{2}\equiv\varphi_{2}(Y^{1},D_{1},\xi^{*})$.
In period $t=2$, a rational individual with learning chooses:
$D_{2}=\operatorname*{argmax}_{d_{2}\in\{0,1\}}\underset{\mu_{2}(d_{2},Y^{1},D_{1},\xi^{*})}{\underbrace{\int u_{2}(d_{2},y_{2})dF_{Y_{2}(D_{1},d_{2})|Y^{1},D_{1},\xi^{*}}(y_{2})}}+\eta_{2}(d_{2}),$ (33)
where $F_{Y_{2}(D_{1},d_{2})|Y^{1},D_{1},\xi^{*}}$ denotes the posterior for $Y_{2}(D_{1},d_{2})$ according to $\kappa_{2}$.111111 Formally, $F_{Y_{2}(D_{1},d)|Y^{1},D_{1},\xi^{*}}(y)=\int\left[\int 1\{\lambda_{2}(D_{1},d,Y^{1},a,e)\leq y\}dF_{\varepsilon_{2}}(e)\right]dF_{\alpha(d)|\kappa_{2}}(a)$, where $F_{\alpha(d)|\kappa_{2}}$ denotes the posterior for $\alpha(d)$ at the start of period $t=2$ given $(Y^{1},D_{1},\xi^{*})$. In period $t=1$, a rational, forward-looking individual chooses:
$D_{1}=\operatorname*{argmax}_{d_{1}\in\{0,1\}}\int\left[u_{1}(d_{1},y_{1})+\rho V_{2}(Y_{0},y_{1},d_{1},\xi^{*})\right]dF_{Y_{1}(d_{1})|Y_{0},\xi^{*}}(y_{1})+\eta_{1}(d_{1}),$ (34)
where $\rho$ is the time discount factor, $F_{Y_{1}(d)|Y_{0},\xi^{*}}$ is the prior of $Y_{1}(d)$ implied by $\kappa_{1}$, and:
$V_{2}(Y_{0},Y_{1},D_{1},\xi^{*})\equiv\int\max_{d_{2}\in\{0,1\}}[\mu_{2}(d_{2},Y_{0},Y_{1},D_{1},\xi^{*})+\tilde{\eta}_{2}(d_{2})]F_{\eta_{2}}(\tilde{\eta}_{2}).$
The solutions of (33) and (34) imply that rational treatment choices by forward-looking individuals with learning necessarily takes the form of (32) with persistent individual fixed effect in the prior and posterior: $\xi_{0}=\xi_{1}(\cdot)=\xi^{*}$. (End of Example 1.)
The next proposition shows, in an observational context where potential outcomes and realized treatments are determined as in (31) and (32), a conditional independence condition is sufficient for sequential exchangeability, but does not imply parallel trends in general.
Proposition 8
Suppose the potential outcomes and realized treatments are determined as in (31) and (32), and:
$(\alpha,\varepsilon_{1},\varepsilon_{2})\perp(\xi_{0},\xi_{1}(d_{1}),\eta_{1},\eta_{2})\mid Y_{0}\text{ for all }d_{1}\text{.}$ (35)
Then sequential exchangeability in Assumption 2b holds while $E[Y_{2}(0,0)-Y_{1}(0)|D_{1},D_{2},Y_{0}]$ is in general non-degenerate in $(D_{1},D_{2})$.
The conditional independence in (35) posits that the fixed effects and idiosyncratic errors influencing posteriors are independent from those determining the potential outcomes, after controlling for the initial outcome $Y_{0}$. It does not further restrict learning.
For example, consider a simplistic case where $\varepsilon^{2},\eta^{2}$ are idiosyncratic noises independent from all other variables, and where $\xi_{1}(d_{1})=\xi_{0}$ almost surely. That is, the evolution of the posterior depends on an initial shifter of the prior $\xi_{0}$ and observed history $(D^{t-1},Y^{t-1})$. Then (35) holds when $\alpha(\cdot)$ is independent from $\xi_{0}$ given $Y_{0}$.
The general failure of the PT condition in Proposition 8 is intuitive. With forward-looking and learning, the treatment choices $D^{2}=(D_{1},D_{2})$ are a non-trivial function of a realized history of observed outcomes $Y_{0},Y_{1}(D_{1})$. Hence conditioning on different realizations of $D^{2}$ amounts to conditioning on different events defined over the sample space of $(Y_{0},\alpha(\cdot),\varepsilon_{1},\xi_{0},\xi_{1}(\cdot),\eta^{2})$. Because $Y_{2}(0,0)-Y_{1}(0)$ is a function of $(Y_{0},\alpha(\cdot),\varepsilon^{2})$, its distribution (and mean) generally vary with the conditioning event defined by $D^{2}$. This non-degeneracy holds in general, even after controlling for the realization in $Y_{0}$.